OpenStudy (anonymous):

Diff equation: xy' = y + 1 My answer is: y = +/- xe^c - 1 is this correct?

OpenStudy (anonymous):

before evaluation the integral i have dy/(y+1) = dx/x

myininaya (myininaya):

subtract y on both sides \[xy'-y=1 \] divide both sides by x \[y'-\frac{1}{x}y=\frac{1}{x}\] integrating factor is \[v=e^{\int\limits_{}^{}\frac{-1}{x} dx}=e^{-\ln(x)} =\frac{1}{x} , x>0\] \[\frac{1}{x}y'-\frac{1}{x^2}y=\frac{1}{x^2}\] now we can write \[(\frac{1}{x}y)'=\frac{1}{x^2}\] now integrate both sides

OpenStudy (amistre64):

yeah, i was looking for a way to get rid of the x klingon lol

myininaya (myininaya):

\[\frac{1}{x}y=\frac{-1}{x}+C\]

OpenStudy (amistre64):

seperation of variables would work as well, yes

myininaya (myininaya):

yes separation of variables works too

OpenStudy (amistre64):

xy' = y + 1 dy/dx x = (y+1) dy/(y+1) x = dx dy/(y+1) = dx/x ; int the whole thing ln|y+1| = ln|x| + C

OpenStudy (anonymous):

so far so good @ amistre64

OpenStudy (amistre64):

e^ both sides then y+1 = e^(ln|x|+C) y+1 = e^(ln|x|) * e^C y+1 = x * C y = Cx - 1 looks about right to me

OpenStudy (amistre64):

e^C is just a constant so we can rewrite it as C

OpenStudy (anonymous):

but is it +/- Cx - 1

OpenStudy (amistre64):

I think the generic constant C makes up for any +- business

OpenStudy (amistre64):

the wolf agrees i think

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=xy%27+%3D+y%2B1

OpenStudy (anonymous):

ahh oke.. but i hadn't simplified it to C so this is correct? y = +/- xe^c - 1 :)

OpenStudy (amistre64):

thats "correct" yes; but i dont think its simplified as much

myininaya (myininaya):

e^c gives us positive constant actually

OpenStudy (amistre64):

it could be more generic that is

myininaya (myininaya):

e^c is never negative

OpenStudy (amistre64):

"e^c gives us a positive constant" he responded in a snitty tone. lol

myininaya (myininaya):

he?

OpenStudy (amistre64):

yes, im a he ;)

myininaya (myininaya):

you're weird

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

ofcourse! e^x is always positive so i can leave the negative sign out :)

OpenStudy (amistre64):

yep, leave of the negative sign

OpenStudy (anonymous):

okee thank you guys =D and yeaahhh i got an answer correctly ^^

myininaya (myininaya):

but why would you want to say plus or minus to begin with?

OpenStudy (amistre64):

prolly just to cover the bases

OpenStudy (anonymous):

because |y+1| were absolute

myininaya (myininaya):

oh i did my a different way

OpenStudy (anonymous):

and i want to have y+1

OpenStudy (anonymous):

thats why i added the plus and minus signs

myininaya (myininaya):

ok